# LLM-Enhanced User-Item Interactions: Leveraging Edge Information for Optimized Recommendations

### v1

@@ Abstract.
The extraordinary performance of large language models has not only reshaped the research landscape in the field of natural language processing but has also demonstrated its exceptional applicative potential in various domains. However, the potential of these models in mining relationships from graph data remains under-explored. Graph neural networks, as a popular research area in recent years, have numerous studies on relationship mining. Yet, current cutting-edge research in graph neural networks has not been effectively integrated with large language models, leading to limited efficiency and capability in graph relationship mining tasks. A primary challenge is the inability of LLMs to deeply exploit the edge information in graphs, which is critical for understanding complex node relationships. This gap limits the potential of LLMs to extract meaningful insights from graph structures, limiting their applicability in more complex graph-based analysis. We focus on how to utilize existing LLMs for mining and understanding relationships in graph data, applying these techniques to recommendation tasks. We propose an innovative framework that combines the strong contextual representation capabilities of LLMs with the relationship extraction and analysis functions of graph neural networks for mining relationships in graph data. Specifically, we design a new prompt construction framework that integrates relational information of graph data into natural language expressions, aiding LLMs in more intuitively grasping the connectivity information within graph data. Additionally, we introduce graph relationship understanding and analysis functions into LLMs to enhance their focus on connectivity information in graph data. By enhancing the understanding of graph relationships by LLMs, our framework provides more comprehensive, accurate, and personalized recommendations. Our evaluation on real-world datasets demonstrates the framework’s ability to understand connectivity information in graph data and to improve the relevance and quality of recommendation results. Our code is released at: https://github.com/anord-wang/LLM4REC.git
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
We aim to explore new methods for applying large language models to recommendation systems, especially when dealing with data containing a large amount of edge information. Our goal is to develop a recommendation system that can effectively utilize this edge information while integrating the powerful text-processing capabilities of LLM. Through this approach, not only can we improve the relevance and accuracy of recommendations, but we can also provide a richer and more personalized experience.
In recent years, significant progress has been made in the development of LLMs such as BERT and GPT. These models demonstrate excellent performance in natural language understanding and generation, showing enormous potential in applications in multiple fields  (Han et al., 2021) (Peng et al., 2023). With the advancement of technology, LLM’s ability to handle complex text data continues to enhance, providing new solutions for various tasks and opening a new chapter in intelligent system research (Birhane et al., 2023). In this context, combining LLM with recommendation systems has become a cutting-edge and revolutionary research field  (Lin et al., 2023). Traditional recommendation systems focus on analyzing users’ behavior data, while when combined with LLM, recommendation systems can understand explicit feedback from users and mine deeper into their implicit needs and preferences  (Ren et al., 2023). This combination provides new possibilities for improving the accuracy and satisfaction of recommendation systems.
A major challenge in current recommendation systems is that although data contain a large amount of edge information (such as the relationship between users and items), this information is not fully utilized in LLMs, especially in its key attention mechanism (Wang et al., 2023). Even though people have used this edge information to construct a variety of prompts  (Wu et al., 2023), existing methods of building recommendation systems using LLMs cannot consider edge information structurally. This raises a research question: how to effectively integrate graph information within the framework of LLMs to improve the performance of recommendation systems?
To overcome the challenges faced by existing recommendation systems in processing data containing a large amount of edge information, we propose an innovative prompt mechanism that can convert the connection relationship between users and items, as well as the background information of items, into natural language text. Our method considers the direct relationship between users and items and enriches the model’s understanding by constructing a second-order relationship between items, a complex association not directly present in traditional recommendation data. At the same time, we also introduce an improved attention mechanism. This mechanism calculates spatial information between nodes in the graph based on the connection relationship between users and items and directly integrates this information into the computation of the attention mechanism as edge encoding. This method enhances the model’s ability to process graph-structured data and greatly improves the accuracy and efficiency of recommendation systems through a deep understanding of the multidimensional relationships between users and items within the model. Through these innovations, we can effectively integrate complex graph structure information into the LLMs framework, greatly improving the model’s comprehensive understanding of users’ behavior and items’ characteristics. This enhances the model’s ability to simulate real-world user-item interactions and lays a solid foundation for generating more accurate and personalized recommendation results. In our model, each recommendation is not only a simple response based on users’ historical behavior but also an intelligent decision-making process that deeply and comprehensively considers users’ preferences and items’ attributes.
To validate the effectiveness of our method, we conduct a series of experiments on different recommendation datasets, including validation experiments on the novel prompt mechanism and improved injection attention mechanism. These experiments demonstrate the performance improvement of our method on recommendation tasks and reveal its potential in handling complex user-item relationships.
Our contributions are as follows:
  • An innovative way of combining LLMs with recommendation systems: By creatively integrating LLMs’ deep text understanding ability with users’ behavior analysis of recommendation systems, our method can more comprehensively understand users’ and items’ information. Our method effectively utilizes the language processing capabilities of LLMs, bringing new dimensions and depth to traditional recommendation logic, thereby ensuring recommendation quality while also increasing the diversity and innovation.
  • A new prompt strategy: We introduce a new prompt mechanism that can transform the relationship between users and items, as well as the background information of items, into natural language form. In addition, by constructing second-order relationships between items, we can uncover deeper correlations between items, thereby providing more comprehensive and detailed recommendations.
  • A novel fusion method for edge information: We propose a new approach that directly embeds the edge information of graph data into the attention mechanism of LLMs. This method effectively utilizes the connection information in the graph structure, enhancing the model’s ability to handle complex user-item interactions.
@@ 2.1. Important Definitions
Prompt. Prompts are designed to guide generative large language models to generate specific responses. They serve as guides for the model to generate text tokens in specific contexts or styles (Liu et al., 2023b).
@@ 2.2. Problem Statement
Consider the existence of $I$ users and $J$ items, let $X_{ij}$ be the binary interaction (e.g., purchase) matrix between user $i$ and item $j$. Besides, we collect user descriptions, item descriptions (e.g., prices, brand, category, title), user reviews for items, explanations of user purchase reasons. We denote $T_{i}$ as the descriptions of the user $i$, $T_{j}$ as the descriptions of the item $j$, $T_{ij}$ as the joint texts of the user $i$ and item $j$, such as user reviews and purchase reasons for items. We unify all textual descriptions into $T$ that includes $N$ sequences, $k$ indexes the tokens in each sequence, and $T_{nk}$ is the $k-th$ token in the $n-th$ sequence. Our goal is to leverage LLMs and graphs to develop a generative recommendation system that takes a prompt including a userID and a user’s historical interaction records with items, and generate product recommendations to the user.
@@ 3. Leveraging LLM and Graphs for Recommender Systems
Figure 1. Overview of Leveraging LLM and Graphs for Recommendation Systems
@@ 3.1. Framework Overview
3.2. Graph Structured Attentive LLM based Generative Recommendation Backbone
@@ 3.2.1. GPT2 as LLM Base Model
Our base model is the GPT-2 (Radford et al., 2019). The original GPT-2 utilizes the Transformer architecture, pre-trained on vast text datasets to predict subsequent words in sequences. With a multi-layer structure containing attention heads, it scales up to billions of parameters for enhanced pattern recognition. The model supports conditional text generation and offers various sampling strategies for generating text. In GPT-2, the attention mechanism is to weigh the importance of different words in a sentence. It operates by computing attention scores for each word in the input sequence based on their relevance to each other. These scores determine how much attention the model should pay to each word when generating the next word in the sequence. By attending to relevant parts of the input text, GPT-2 can capture long-range dependencies and generate coherent and contextually relevant output.
@@ 3.2.2. Integrating Two Structure Knowledge for Graph Attentive LLM
When performing generative recommendations, we obtain recommendation results in the form of text generation to connect items to users. As a result, users and items are usually regarded as tokens in a text sequence for pre-training. In the real world, users interact (e.g., rate, purchase) with items. Such interactions can be modeled as a graph where users or items are seen as nodes, and user-user, item-item, user-item connectivity is seen as edge weights, representing a kind of graph-structured information propagation-driven edge knowledge. In other words, user and item tokens are not just simply independent entities in a sequence. The LLM should not just learn user and item embeddings by paying attention to their mutual relevance in a sequence. It is critical to leverage the graph-structured edge knowledge to improve LLM for recommendations.
Firstly, we propose a graph structure knowledge attentive LLM method in order to integrate graph knowledge into recommendation systems. Specifically, we incorporate the edge connectivity between users and items into attention weight calculation (Gao et al., 2020; Ying et al., [n. d.]). Our idea is to leverage Graph Neural Networks (GNNs) to describe the relationships between nodes by modeling their connectivity (direct relationship) and spatial information (indirect relationship) in the graph. Formally, the edge information, denoted by the $R$-term is used to calculate the attention weights in the graph-structured attention mechanism, which is given by:
@@
where $Q$, $K$, and $V$ are queries, keys, and values respectively, while $R$ represents the relationship encoding extracted from graph knowledge. The $\sqrt{d_{k}}$ is the dimensionality (size) of the key vector to enforce a normalization effect.
Secondly, we identify two kinds of important graph structural knowledge: 1) the direct (first-order) connectivity and 2) the indirect (high-order path) connectivity, among users and items. Correspondingly, the graph structured relational attention $R$ term is composed of two distinct parts of the graph topology:
@@
(3) $R_{ij}^{conn}=\begin{cases}1,&\text{if there is a direct connection between node }i\text{ and }j\\ 0,&\text{otherwise}\end{cases}.$
In this matrix, 1 indicates there is an edge between two nodes, and 0 indicates there is no edge between two nodes. The second part, $R_{\text{path}}$, represents a normalized shortest path score between nodes, which is computed based on the entire graph. The shortest path information is essential because it reflects the indirect relationships of node pairs and the degrees of separation or distance between nodes, which can be highly informative for understanding complex graph structures. The normalized shortest path score is calculated using the shortest path matrix $P$, where each element $P_{ij}$ is defined as the minimum path length among all possible paths from node $i$ to node $j$. The $P_{ij}$ is given by: $P_{ij}=\text{min}\{\text{path length}|\text{all paths from node }i\text{ to node }j\}.$ Later, we introduce a damping factor $\delta$ in order to adjust the influence of distant nodes. This is achieved by inverting and normalizing the path lengths in the matrix. The modified shortest path matrix $R^{\text{path}}$ is defined as:

### v2

@@ Abstract.
Graph recommendation methods, representing a connected interaction perspective, reformulate user-item interactions as graphs to leverage graph structure and topology to recommend and have proved practical effectiveness at scale. Large language models, representing a textual generative perspective, excel at modeling user languages, understanding behavioral contexts, capturing user-item semantic relationships, analyzing textual sentiments, and generating coherent and contextually relevant texts as recommendations. However, there is a gap between the connected graph perspective and the text generation perspective as the task formulations are different. A research question arises: how can we effectively integrate the two perspectives for more personalized recsys? To fill this gap, we propose to incorporate graph-edge information into LLMs via prompt and attention innovations. We reformulate recommendations as a probabilistic generative problem using prompts. We develop a framework to incorporate graph edge information from the prompt and attention mechanisms for graph-structured LLM recommendations. We develop a new prompt design that brings in both first-order and second-order graph relationships; we devise an improved LLM attention mechanism to embed direct the spatial and connectivity information of edges. Our evaluation of real-world datasets demonstrates the framework’s ability to understand connectivity information in graph data and to improve the relevance and quality of recommendation results. Our code is released at: https://github.com/anord-wang/LLM4REC.git.
Large Language Models, Recommender System, Attention Mechanism, Graph
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 1publicationmonth: 8ccs: Information systems Data miningccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recommender Systems (RecSys) analyzes multi-source user behavioral data and models complex user-item interactions for accurate and personalized recommendations. In networked data, graphs play a role in RecSys due to their ability to represent user-item relationships and capture complex user-user and item-item connections. Example methods include graph neural networks, random walks, matrix factorization, and PageRank  (Brin and Page, 1998). In textual data, Large Language Models (LLMs) have demonstrated amazing abilities to model user languages, understand behavioral contexts, capture user-item semantic relationships, analyze textual sentiments, and generate coherent and contextually relevant texts as recommendations. In this paper, we aim to integrate graphs (a connected interaction aspect) and LLMs (a textual generative aspect) to improve the relevance, accuracy, and personalization of RecSys.
In prior literature, graph-based methods, such as random-walk  (Page et al., [n. d.]), graph neural networks  (Kipf and Welling, 2016), graph factorization  (Ahmed et al., 2013), and graph contrastive learning  (You et al., 2020) have reformulated RecSys as a task of link prediction or graph classification and have demonstrated the effectiveness of such connected graph perspective. Recently, LLMs have seen remarkable advancements. The Transformer Architecture and Attention Mechanisms (Vaswani et al., 2017) was introduced in 2017, followed by large-scale pre-trained language models like BERT  (Devlin et al., 2018). In 2018, OpenAI introduced the GPT (Generative Pre-trained Transformer) model  (Radford et al., 2018), which is one of the first large-scale language models based on the Transformer Encoder Architecture. Subsequent improved versions, such as GPT-2 (2019)  (Radford et al., 2019), GPT-3 (2020)  (Brown et al., 2020), GPT-4 (2023)  (Achiam et al., 2023), and GPT-4o (2024), increased model size, leading to improved performance and smarter capabilities like few-shot learning and open-ended text generation. Researchers found that the LLMs, with their ability to capture rich contextual representations and handle long-range dependencies, showed potential in understanding user preferences and item characterizations. Geng et al. proposed the concept of Recommendation as Language Processing in 2022 and leveraged pre-trained LLMs like T5 (Raffel et al., 2020) to perform sequential recommendation, rating prediction, explanation generation, and direct recommendation.  (Geng et al., 2022) Thereafter, there are many studies that integrate LLM into RecSys by improving prompt engineering, fairness, and debiasing, efficient model architectures, explainability, ethical issues under RecSys  (Zhao, 2022; Sun et al., 2019; Hou et al., 2022; Liu et al., 2023a; Dai et al., 2023).
However, there is a clear gap between LLMs-based RecSys and graph-based RecSys. In the LLMs setting, RecSys is seen as an auto-regressive textual generation task, while in the graph setting, RecSys is seen as a link prediction task. The two perspectives are fundamentally different. A major challenge is how we can effectively integrate the connected graph perspective (edge information) into LLMs for effective RecSys.
Our Insights: Incorporating Graph Edge Information into Large Language RecSys via Prompt and Attention Innovations. To fill the gap between LLMs-based RecSys and graph based RecSys, we renovate prompt design and attention mechanism. Firstly, we develop a prompt design to take into account not just user profiles and item descriptions, but also user-item edge interactions that particularly include both direct relationships between users and items and second-order relationships between items, a complex association not directly present in traditional recommendation data. Such graph edge-enhanced prompt design equips LLMs with better abilities of contextual understanding of user-item relationships. Secondly, we introduce a graph structure knowledge attention mechanism into the pre-trained transformer model. The attention mechanism leverages not just the relationships between nodes by modeling their connectivity (direct relationship) but also spatial information (indirect relationship) in the graph.
Solution Summary. We regard user and item nodes as target recommendation tokens and reformulate recommendations as a probabilistic generative problem using prompts. Specifically, we first construct a user-item interaction graph and develop a graph structure-guided attentive LLM backbone with a new neural attentive decoder to model these connections. The training is achieved by creating text prompts for all users and items, including interaction events and crowd contextual prompts, using these prompts to pre-train the backbone by maximizing text generative likelihood to learn contextual knowledge relevant to recommendations. Besides, we convert a user’s interaction history into past-tense texts and combine them with future-tense triggers to fine-tune the LLM backbone, by minimizing recommendation errors. Finally, we leverage the fine-tuned graph attentive LLM backbone to make recommendations based on the personalized rating or purchase history and predictive triggers of test users.
Our Contributions. 1) Framework. We tackle a graph-structured LLM RecSys task and develop a framework to incorporate graph edge information from the prompt and the attention mechanism. 2) Computing. We develop a new prompt design that brings in both first-order and second-order graph relationships; we devise an improved LLM attention mechanism to embed direct the spatial and connectivity information of edges. 3) Experiments. We present extensive experiments on a series of recommendation datasets to demonstrate the performance of recommendation tasks and the effectiveness of graph-structured prompt and attention mechanisms.
@@ 2. Preliminaries
2.1. Symbol Definition
Table 1 summarizes the key symbols used throughout this paper, covering user-item interactions, textual representations, model parameters, and graph structures. It provides a reference for understanding the mathematical formulations and concepts in our proposed approach.
Table 1. List of Symbols Used in This Paper.
Symbol Description
$I$ The number of users in the dataset.
$J$ The number of items in the dataset.
$X_{ij}$ The binary interaction between user $i$ and item $j$.
$T_{i}$ The descriptions of the user $i$.
$T_{j}$ The descriptions of the item $j$.
$T_{ij}$ The joint texts of the user $i$ and item $j$.
$T$ The textual descriptions.
$N$ The number of sequences in $T$.
$T_{nk}$ The $k\textit{-}th$ token in the $n\textit{-}th$ sequence.
$Q$ The query vector in the Transformer structure.
$K$ The key vector in the Transformer structure.
$V$ The value vector in the Transformer structure.
$\sqrt{d_{k}}$ The dimensionality (size) of the key vector to enforce a normalization effect.
$R$ The relationship encoding extracted from graph knowledge.
$R^{\text{conn}}$ The direct connection relationships between nodes.
$R^{\text{path}}$ The normalized shortest path score between nodes.
$P$ The shortest path matrix.
$P_{ij}$ The minimum path length among all possible paths from node $i$ to node $j$.
$\delta$ The normalization factor between 0 and 1.
$\mathcal{L}_{\text{pre-train}}$ The objective function for maximizing text generation likelihood during pre-training.
$t_{i}$ The tokens in the next token prediction task.
$\mathcal{L}_{\text{fine-tune}}$ The objective function for optimizing recommendation accuracy during fine-tuning.
$\mathcal{S}_{i}$ The list of recommended items $\mathcal{S}_{i}$ to user $i$.
$\text{Pr}_{i}$ The personalized predictive prompt for user $i$ in fine-tuning.
$\Theta$ The set of LLM parameters.
@@ 2.2. Important Definitions
Prompt. Prompts are designed to guide generative LLMs to generate specific responses. They serve as guides for the model to generate text tokens in specific contexts or styles (Liu et al., 2023b).
@@ 2.3. Problem Statement
Consider the existence of $I$ users and $J$ items, let $X_{ij}$ be the binary interaction (e.g., purchase) matrix between user $i$ and item $j$. Besides, we collect user descriptions, item descriptions (e.g., prices, brand, category, title), user reviews for items, and explanations of user purchase reasons. We denote $T_{i}$ as the descriptions of the user $i$, $T_{j}$ as the descriptions of the item $j$, $T_{ij}$ as the joint texts of the user $i$ and item $j$, such as user reviews and purchase reasons for items. We unify all textual descriptions into $T$ that includes $N$ sequences, $k$ indexes the tokens in each sequence, and $T_{nk}$ is the $k\textit{-}th$ token in the $n\textit{-}th$ sequence. Our goal is to leverage LLMs and graphs to develop a generative recommender system that takes a prompt, including a user ID and a user’s historical interaction records with items, and generates product recommendations to the user.
@@ 3. Leveraging LLM and Graphs for Recommender Systems
Figure 1. Overview of the proposed graph-attentive LLM-based recommender system. The framework consists of four key steps: (1) Developing a graph-attentive LLM backbone, which integrates multi-source textual information and user-item interaction graphs to enhance representation learning; (2) Pre-training with crowd contextual prompts, where structured text prompts encode user and item descriptions along with interaction facts to establish contextual knowledge; (3) Fine-tuning with personalized predictive prompts, converting past user interactions into structured prompts with future prediction triggers to optimize recommendation accuracy; and (4) Generative recommendation, where the fine-tuned LLM backbone predicts items based on user history and interaction patterns.
@@ 3.1. Framework Overview
Given user descriptions, item descriptions, user textual reviews for items, and the user-item interaction (e.g., purchase, rating) graph, we aim to leverage and connect generative LLM, textual generation, and user-item interaction graphs to advance recommender systems (Wu et al., 2023). Fig. 1 shows the four major steps for building our recommender system: 1) developing graph attentive LLM recommendation backbone; 2) pre-training the backbone with crowd contextual prompts; 3) fine-tuning of the backbone with personalized predictive prompts; 4) generative recommendation.
In Step 1, considering the existence of multi-source textual information, including user descriptions, item descriptions, and user reviews for items, we propose to leverage LLM to learn the generation of these texts to model the representations of user preferences and item functionalities. Aside from texts, the user-item interaction graph can provide two types of signals: i) user nodes and item nodes as tokens and ii) graph structures about the direct (first-order) connectivity among users and items and the indirect (second-order) connectivity among items, users, or user-item pairs. To leverage the user and item tokens, we add the user and item tokens to enrich the texts; to leverage the graph structures, we develop a graph knowledge-guided attentive LLM backbone, particularly with a new neural attentive decoder structure, to model the first-order and second-order connectivity graph knowledge. Our graph-attentive LLM backbone reformulates recommendations into a probabilistic generative problem in response to prompts.
Step 2 is to pretrain the graph-attentive LLM backbone. Specifically, we first construct the text prompts of all users and items, including the texts of user descriptions, item descriptions, and user reviews for items, the fact or event checking texts of whether a user interacts with (e.g., rates or purchases) items, as crowd contextual prompts. We utilize the crowd contextual prompts to pre-train the backbone by maximizing text generative likelihood. So, LLM can gain contextual knowledge of a world of recommendation.
@@
Finally, given a test user with the corresponding personalized rating or purchase history and a predictive trigger, Step 4 leverages the fine-tuned graph attentive LLM backbone to recommend items to the test user.
3.2. Graph-Structured Attentive LLM-Based Generative Recommendation Backbone
@@ 3.2.1. GPT2 as LLM Base Model
Our base model is the GPT-2 (Radford et al., 2019). The original GPT-2 utilizes the Transformer architecture, pre-trained on vast text datasets to predict subsequent words in sequences. A multi-layer structure containing attention heads scales up to billions of parameters for enhanced pattern recognition. The model supports conditional text generation and offers various sampling strategies for generating text. In GPT-2, the attention mechanism is to weigh the importance of different words in a sentence. It operates by computing attention scores for each word in the input sequence based on their relevance to each other. These scores determine how much attention the model should pay to each word when generating the next word in the sequence. By attending to relevant parts of the input text, GPT-2 can capture long-range dependencies and generate coherent and contextually relevant output.
@@ 3.2.2. Integrating Two Structure Knowledge for Graph Attentive LLM
When performing generative recommendations, we obtain recommendation results in the form of text generation to connect items to users. As a result, users and items are usually regarded as tokens in a text sequence for pre-training. In the real world, users interact (e.g., rate, purchase) with items. Such interactions can be modeled as a graph where users or items are seen as nodes, and user-user, item-item, and user-item connectivity is seen as edge weights, representing a kind of graph-structured information propagation-driven edge knowledge. In other words, user and item tokens are not simply independent entities in a sequence. The LLM should not just learn user and item embeddings by paying attention to their mutual relevance in a sequence. It is critical to leverage the graph-structured edge knowledge to improve LLM for recommendations.
Firstly, we propose a graph structure knowledge-attentive LLM method to integrate graph knowledge into recommender systems. Specifically, we incorporate the edge connectivity between users and items into attention weight calculation (Gao et al., 2020). Inspired by the Graphormer  (Ying et al., [n. d.]), we leverage Graph Neural Networks (GNNs) to describe the relationships between nodes by modeling their connectivity (direct relationship) and spatial information (indirect relationship) in the graph. Formally, the edge information, denoted by the $R$-term, is used to calculate the attention weights in the graph-structured attention mechanism, which is given by:
@@
where $Q$, $K$, and $V$ are queries, keys, and values respectively, while $R$ represents the relationship encoding extracted from graph knowledge. The $\sqrt{d_{k}}$ is the dimensionality (size) of the key vector to enforce a normalization effect. The structural scores $R$ are added to the original attention scores because there should be additional scores between connected nodes. It means that connected nodes should pay more attention to each other. This is a way to reflect graph edges in the attention mechanism.
Secondly, we identify two kinds of important graph structural knowledge: 1) the direct (first-order) connectivity and 2) the indirect (high-order path) connectivity, among users and items. Correspondingly, the graph-structured relational attention $R$ term is composed of two distinct parts of the graph topology:
@@
(3) $R_{ij}^{\text{conn}}=\begin{cases}1,&\text{if there is a direct connection between node }i\text{ and }j\\ 0,&\text{otherwise}\end{cases}.$
In this matrix, 1 indicates there is an edge between two nodes, and 0 indicates there is no edge between two nodes. The second part, $R^{\text{path}}$, represents a normalized shortest path score between nodes, which is computed based on the entire graph. The shortest path information is essential because it reflects the indirect relationships of node pairs and the degrees of separation or distance between nodes, which can be highly informative for understanding complex graph structures. The normalized shortest path score is calculated using the shortest path matrix $P$, where each element $P_{ij}$ is defined as the minimum path length among all possible paths from node $i$ to node $j$. The $P_{ij}$ is given by: $P_{ij}=\text{min}\{\text{path length}|\text{all paths from node }i\text{ to node }j\}.$ Later, we introduce a damping factor $\delta$ to adjust the influence of distant nodes. This is achieved by inverting and normalizing the path lengths in the matrix. The modified shortest path matrix $R^{\text{path}}$ is defined as: